Suppose that 55% of all babies born in a particular hospital are boys. If 7 babies born in the hospital are randomly selected, what is the probability that at least 3 of them are boys?

Carry your intermediate computations to at least four decimal places, and round your answer to two decimal places.
Answer #1

Solution

Given that ,

p = 0.55

q = 1 - p = 1 - 0.55 = 0.45

n = 7

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X 3 ) = 1 - P( x <3)

= 1 - P(X = 0) P(X = 1) P(X = 2)

= 1 - (7 C 0) * 0.55 0 * (0.45)7  (7 C 1) * 0.55 1 * (0.45)6  (7 C 2) * 0.55 2 * (0.45)5

=1-0.1529

probability=0.8471

answer=0.85